A uniform electric field of 8 V/m is directed in the negative x direction. If one moves a distance of 2.7 meters in the positive x direction, how much does the electric potential change in Volts?
1
Expert's answer
2013-02-19T07:41:54-0500
If you want to find the change of electric potential along x axis, you have to multiply difference of distances, that is 2.7 meters, by the electric field itself, if it is uniform, that is 8 V/m. Therefore we get

V = x*E =2.7 m *8 Volts/m = 21.6 Volts